find the area of the quardilateral whose coordinates are (0,0),(3,0),(0,3),(7,5)
Answer by Alan3354(69443)

find the area of the quardilateral whose coordinates are (0,0),(3,0),(0,3),(7,5)
-----------
Label the points A to D in order (0,0),(3,0),(7,5),(0,3)
A B C D A
0 3 7 0 0
0 0 5 3 0
-------------
Add the upward diagonal products:
0*3 + 0*7 + 5*0 + 3*0 = 0
Add the downward diagonal products:
0*0 + 3*5 + 7*3 + 0*0 = 36
-------------
Area = (36 - 0)/2
Area = 18 sq units.